# TreeEval: Benchmark-Free Evaluation of Large Language Models through Tree Planning

### v1

@@ 4.2 Performance of TreeEval
We display the performance of TreeEval in Table 1, from which we have the following observations: (1) Among all the comparable evaluation methods, our method is able to achieve the highest correlation coefficient with the rankings of AlpacaEval2.0 on the indicators of both $\rho$ and $\tau$. AlpacaEval2.0 is commonly viewed as the recognized LLM evaluation leader-board and the high consistency between our ranks indicates the reliability of our method. (2) Our method is able to complete the evaluation procedure with only $45$ questions in average while the other evaluation methods require much more questions to generate an evaluation result. This indicates that our evaluation is efficient on evaluating LLMs with minimum questions. (3) Since we treat Mistral-7B-Instruct-v0.2 as the reference for the pairwise comparison, we notice the larger gap between the evaluated LLM and the reference is, the less test questions are proposed in the evaluation session, which shows that tree planning indeed meets our expected motivation. We further display the pairwise correlation in Appendix  8.5 to show the correlation between TreeEval and AlpacaEval is also high.
Figure 5: Re-run TreeEval 5 times for various LLMs.
4.3 Further Analysis
We conduct further analysis to verify the effect of TreeEval [^4] More pairwise comparison for different model pairs and method pairs analysis can be found in Appendix 8.4,8.5..
Fine-grained Evaluation. From the Figure 5, we can observe that different LLMs excel in various knowledge domains. The performance of the same model may vary across different fields. For example, Yi-34B-Chat is short in Travel and Shopping while it demonstrates relatively good performance on other topics. Through our TreeEval, we can diagnose an LLM with the fine-grained results in diverse domains.
Robustness of TreeEval. We draw Figure 5 to show the evaluation results of different LLMs in multiple runs. We can see that, for a given LLM under evaluation, conducting multiple repeated experiments yields relatively similar scores when the examiner’s temperature is set to 1. The low variance indicates that TreeEval is able to generate stable and robust results.
@@ 4.3 Further Analysis
Case Studies. In Figure 3, it’s clear that TreeEval effectively identifies performance gaps between LLMs. For instance, Mistral-7B-Instruct-v0.2 notably outperforms WizardLM-13B-V1.2, reflected in a smaller tree. Conversely, when models perform similarly, like Mistral-7B-Instruct-v0.2 and Xwin-LM-13B-V0.1, TreeEval constructs a larger tree to discern subtle performance differences.
5 Conclusions
@@
6 Ethical Considerations
@@
7 Limitations
While we utilize the most powerful general LLM, GPT4, as our examiner in the evaluation process, it is important to recognize that even GPT4 has its limitations and areas where it may not excel. One inherent flaw in our approach is that it may fail when evaluating content that the examiner isn’t proficient in. One potential solution could involve providing more contextual cues to guide the examiner when assessing challenging content. Looking ahead, training a specialized examiner to extract questions from a document repository and evaluate an LLM’s comprehension of the knowledge within that repository could address the issue of evaluating content that the examiner isn’t proficient in.

### v2

@@ Introduction
We display the performance of TreeEval in Table 1, from which we have the following observations: (1) Among all the comparable evaluation methods, our method is able to achieve the highest correlation coefficient with the rankings of AlpacaEval2.0 on the indicators of both $\rho$ and $\tau$. AlpacaEval2.0 is commonly viewed as the recognized LLM evaluation leader-board and the high consistency between our ranks indicates the reliability of our method. (2) Our method is able to complete the evaluation procedure with only $45$ questions in average while the other evaluation methods require much more questions to generate an evaluation result. This indicates that our evaluation is efficient on evaluating LLMs with minimum questions. (3) Since we treat Mistral-7B-Instruct-v0.2 as the reference for the pairwise comparison, we notice the larger gap between the evaluated LLM and the reference is, the less test questions are proposed in the evaluation session, which shows that tree planning indeed meets our expected motivation. We further display the pairwise correlation in (Appendix More analysis) to show the correlation between TreeEval and AlpacaEval is also high.
Further Analysis
We further analyze to verify TreeEval’s effect.444 Due to the page limit, more analyses (i.e., Fine-grained Evaluation and robustness of TreeEval) are provided in (Appendix Fine-grained Evaluation) and (Appendix Robustness of TreeEval), respectively.
LLMs MMLU BBH MT-bench AlpacaEval2.0 TreeEval(Ours)
Acc Acc score Win-Rate #Q Score(Var)
Yi-34B-Chat $73.5$ $71.7$ $8.65$ $27.1$- $2.50(0.000)$
Qwen1.5-110B-Chat $80.4$ $74.8$ $8.88$ $33.7$ $42.67$ $4.03(0.110)$
Meta-Llama-3-70B-Instruct $82.0$ $81.3$ $8.92$ $34.4$ $36.33$ $3.82(0.128)$
Qwen1.5-72B-Chat $75.6$ $65.5$ $8.61$ $36.6$ $31.33$ $3.45(0.089)$
Mixtral-8x7B-Instruct-v0.1 $70.6$ $57.3$ $8.30$ $23.7$ $44.67$ $2.02(0.027)$
vicuna-33b-v1.3 $59.2$ $52.0$ $7.12$ $12.7$ $21.33$ $0.35(0.033)$
$\rho$uparrow $0.82$ $0.71$ $0.71$ $1.0$- $\mathbf{0.94}$
$\tau$uparrow $0.73$ $0.60$ $0.60$ $1.0$- $\mathbf{0.86}$
Table 2: Results of More Powerful Models. Given the strong capabilities of the evaluated models, we use Yi-34B-Chat as the baseline and exclude the AlpacaEval benchmark, which is relatively simple for these models.
More powerful models. To demonstrate the performance of our approach in comparison with more powerful models, Table 2 presents the results of using Yi-34B-Chat as the baseline. We selected some of the most advanced open-source models currently available for testing against Yi-34B-Chat. Our TreeEval achieved performance closest to that of AlpacaEval2. More model results can be found in the (appendix More model results).
Model Yi-34B -Chat Xwin-LM -13B-V0.1 Mistral-7B -Instruct-v0.2 vicuna -33b-v1.3 WizardLM -13B-V1.2 zephyr -7b-beta
Yi-34B-Chat -- $1.88(0.400)$ $1.52(0.010)$ $2.1(0.070)$ $1.21(0.076)$ $1.75(0.143)$
Xwin-LM-13B-V0.1 $3.12(0.400)$- $2.33(0.000)$ $1.53(0.403)$ $1.57(0.109)$ $2.41(0.000)$
Mistral-7B-Instruct-v0.2 $3.48(0.010)$ $2.67(0.000)$- $1.61(0.044)$ $1.10(0.070)$ $2.19(0.003)$
vicuna-33b-v1.3 $2.9(0.070)$ $3.47(0.403)$ $3.39(0.044)$- $2.01(0.374)$ $3.7(0.071)$
WizardLM-13B-V1.2 $3.79(0.076)$ $3.43(0.109)$ $3.90(0.070)$ $2.99(0.374)$- $3.94(0.069)$
zephyr-7b-beta $3.25(0.143)$ $2.59(0.000)$ $2.81(0.003)$ $1.3(0.071)$ $1.06(0.069)$-
Table 3: Our result for each model pairs. The elements in this table represent the scores obtained by comparing models using treeEval, with the column model being compared against the row model.
Pairwise Comparison for different model pairs. We iteratively change the references for the pairwise comparison and the results are shown in Table 3. Choosing the right baseline model is a critical step in our evaluation strategy. We selected the Mistral-7B-Instruct-v0.2 as our baseline, emphasizing the significance of selecting a baseline that accurately reflects the broad insights from pairwise model comparisons. Ideally, a baseline model should have a performance level that is neither too high nor too low, ensuring fair and balanced comparisons across all models. Interestingly, our observations indicate that even a baseline model chosen at random can lead to rankings that closely resemble those from a thorough pairwise evaluation. Thus, it’s feasible to start with a randomly chosen baseline model to set up an initial order of performance. This preliminary order can be refined effectively using the bubble sort method. Given the initial order’s similarity to the final ranking, this refinement process tends towards an $O(n)$ complexity, significantly enhancing the evaluation’s precision and efficiency.
@@
Case studies are presented in (Appendix Case Studies).
Conclusions
@@
Limitations
Using LLMs like GPT-4 as judges introduces potential data leakage risks due to biases in their pre-training data. This can be mitigated by selecting neutral evaluators independent of the assessed models’ training data or randomly rotating evaluators to reduce bias.
While GPT-4 is a powerful examiner, it has limitations, particularly in areas outside its expertise. This can be addressed by providing more contextual guidance during evaluations. In the future, training specialized evaluators to extract questions from document repositories and assess comprehension could ensure more accurate, domain-specific evaluations.
Ethical Considerations
@@
Acknowledgements
The authors would like to thank the anonymous reviewers for their insightful comments. This work is supported by the National Key Research & Develop Plan (2023YFF0725100) and Young Scientists Project of National Natural Science Foundation (Project No. 62206097).
Appendix
More analysis
Figure 4: Re-run TreeEval 5 times for various LLMs.
Fine-grained Evaluation.
From the Figure 4, we can observe that different LLMs excel in various knowledge domains. The performance of the same model may vary across different fields. For example, Yi-34B-Chat is short in Travel and Shopping while it demonstrates relatively good performance on other topics. Through our TreeEval, we can diagnose an LLM with the fine-grained results in diverse domains.
Robustness of TreeEval.
We draw Figure 4 to show the evaluation results of different LLMs in multiple runs. We can see that, for a given LLM under evaluation, conducting multiple repeated experiments yields relatively similar scores when the examiner’s temperature is set to 1. The low variance indicates that TreeEval is able to generate stable and robust results.
Case Studies.
In Figure 5, it’s clear that TreeEval effectively identifies performance gaps between LLMs. For instance, Mistral-7B-Instruct-v0.2 notably outperforms WizardLM-13B-V1.2, reflected in a smaller tree. Conversely, when models perform similarly, like Mistral-7B-Instruct-v0.2 and Xwin-LM-13B-V0.1, TreeEval constructs a larger tree to discern subtle performance differences.
LLMs MMLU BBH AlpacaEval MT-bench AlpacaEval2.0 TreeEval(Ours)
Acc Acc Win-Rate score Win-Rate #Q Score
Mistral-7B-Instruct-v0.2 $70.6$ $46.4$ $92.78$ $8.30$ $14.72$- $2.50$
Vicuna-7b-v1.3 $47.1$ $37.5$ $76.84$ $6.00$ $4.64$ $34.67$ $0.92$
Vicuna-7b-v1.5 $49.9$ $43.3$ $74.50$ $6.17$ $4.79$ $36.33$ $0.96$
Vicuna-13b-v1.3 $52.1$ $43.5$ $82.11$ $6.39$ $7.13$ $38.33$ $1.28$
Vicuna-13b-v1.5 $55.8$ $51.5$ $81.07$ $6.57$ $6.72$ $34.33$ $1.06$
chatglm2-6b $45.5$ $32.7$ $47.12$ $4.96$ $2.76$ $39.67$ $1.12$
alpaca-13b $55.1$ $43.3$ $78.93$ $6.53$ $7.43$ $55.00$ $2.19$
Qwen-14B-chat $66.5$ $58.0$ $79.83$ $6.96$ $7.50$ $36.67$ $1.40$
Starling-LM-7B-alpha $63.9$ $51.5$ $91.99$ $8.09$ $14.24$ $52.00$ $2.36$
$\rho$uparrow $0.57$ $0.21$ $0.25$ $0.31$ $1.0$- $\mathbf{0.70}$
$\tau$uparrow $0.53$ $0.14$ $0.20$ $0.36$ $1.0$- $\mathbf{0.58}$
Table 5: More model results
More model results.
To further demonstrate the performance of our method on different models to be tested, we also show the performance of the vicuna series models as well as chatglm2-6b (Zeng et al. 2022) and alpaca-13b (Chen et al. 2023) in Table 5. But the main analysis still focuses on the previous six models.
Eval Controller Example
For the first aspect: the generated follow-up question should be closely linked to its previous question via topics. For example, in Figure 2, we generate $Q_{2}$ as ”How does 5G technology improve Internet of Things (IoT) applications and smart city initiatives?” Based on the $C_{6}$ we obtained we generated a new question $Q_{6}$ ”In what ways does 5G technology enable advancements in smart home devices and automation?”, $Q_{6}$ is an expansion and extension of $Q_{2}$ on $C_{6}$ ”applications”.
For the second aspect: the generated questions should not be repeated in the existing questions. For example, in Figure 2, $Q_{5}$ was initially ”What are the ethical considerations when designing AI systems for human-machine interaction?” However, this significantly overlapped with our previously established $Q_{3}$ ”What are the key ethical considerations when developing AI technologies for communication platforms?” Consequently, we opted for $Q_{5}$ to be ”What role does AI play in the development of voice-activated systems, and how does it change human-machine interaction?” to ensure variety and specificity in our discussion topics.
Figure 6: To facilitate analysis, Figure 2 is re-placed here.
Prompt for Examiner
I want you to assume the role of the expert and ask a question that expands and reflects your understanding of {topic}. Your task is to ask a question about {topic}. Only through a profound understanding of {topic} can one correctly answer this question. Please adhere strictly to the following 4 task guidelines:
  1. 1.
    Your question should begin with a question word, such as ”what”, ”which”, ”when”, ”where”, ”how”, ”why”, etc.
  2. 2.
    The objective of your question should be to manifest the respondent’s understanding of {topic} and to differentiate respondents based on their comprehension level.
  3. 3.
    Questions should be self-explanatory, not requiring additional context or clarification.
  4. 4.
    Please format your question in the JSON structure provided below. Remember, only output the content in the following format, and nothing else: {{”question”: your question}}
Prompt for Judge
You are assessing two submitted responses to a user’s query based on specific criteria. Evaluate the quality, relevance, accuracy, clarity, and any other relevant factors to determine which response is superior, or if they are equally valuable or lacking. Here is the data for your assessment:
[Query]: {question}
[Response 1]: {answer 1}
[Response 2]: {answer 2}
Assessment Criteria:
  1. 1.
    Relevance to the query: Does the response directly address the user’s question or concern?
  2. 2.
    Accuracy of information: Are the facts or solutions provided in the response correct and reliable?
  3. 3.
    Clarity and comprehensibility: Is the response easy to understand, well-structured, and free of jargon or ambiguity?
  4. 4.
    Completeness: Does the response cover all aspects of the query or offer a comprehensive solution?
  5. 5.
    Additional value: Does the response provide extra insights, tips, or information that enhances the user’s understanding or solves the problem more effectively?
Instructions for Assessment:
  1. 1.
    Identify and focus on the criteria that significantly distinguish the two responses. Disregard criteria that do not offer a clear distinction.
  2. 2.
    Consider any specific aspects of the query and the responses that may require additional factors for a fair comparison. Mention these factors explicitly.
  3. 3.
    Conclude your assessment by deciding which response is better, or if they are tied. Your decision must be based on a coherent evaluation across the mentioned criteria and any additional factors you’ve identified.
Please return your final decision in the following JSON format: {”Eval_result”: ”Response 1”/”Response 2”/”Tie”}
Note: Remember, the output should only contain the decision in the specified JSON format and nothing else.
Prompt for NER
You are asking questions and answers based on a topic you know and based on this topic. Please extract some subtopics from the answers. Here’s an example:
Here is the data:
[Input data]
[topic]: programming languages
[question]: Which programming languages can you write code in?
[answer]: I know python, C++, R language, etc.
[Output Data]
[subtopic] : [”python”,”C++”,”R language”]
now the official question
Here is the data:
[Input data]
[topic]:0
[question]:1
[answer]:2
[Output Data]
Please return your final decision in list format. Remember, you only need to output the content in the following List format, with each element as a subtopic and nothing else. Remember, you only need to output the three most important subtopics in the following List format.
[subtopic] : [”subtopic1”,”subtopic2”,”subtopic3”]